OpenStudy (anonymous):

a 25mL solution of H2SO4 is neutralized by 18 mL of 1.0M of NaOH. What is the concentration of the H2SO4?

OpenStudy (anonymous):

=0.72

OpenStudy (anonymous):

=(18÷1000)×1.0)÷(25÷1000)=(18÷1000)÷(25÷1000)=(18÷1000)×(1000÷25)=18÷25=0.72

OpenStudy (jfraser):

most important is to have the balanced reaction, because you must have the \(mole \space ratio\) first

OpenStudy (jfraser):

\[H_2SO_4 + 2NaOH \rightarrow Na_2SO_4 + 2H_2O\]